# AI-Generated Video Detection via Perceptual Straightening

### v1

@@ 7 Benchmark results
We evaluate ReStraV (with MLP from Section 6) under three settings: A) Vs. SoTA image-based detectors on VidProM [3]; B) Vs. SoTA video detectors on VidProM [3] (evaluating “seen”, “unseen”, and “future” generator scenarios) and on the GenVidBench [17] dataset (in “cross-source”/“generator” (M) scenarios and its “Plants” hard task subclass (P)); C) Vs. a Multi-modal Large Language Model (MLLM) performance (Gemini 1.5 Pro [61]) on the Physics-IQ dataset [26]. The best scores are bold, second best are underlined and the best method overall is highlighted in blue.
7.1 ReStraV vs. image-based detectors.
We evaluate our method ReStraV against eight SoTA image based detectors in Table 2. ReStraV is trained with data and processing from Section 6. We use a balanced test set (40,000 real videos; 10,000 AI-generated for each of four models: Pika [44], VideoCraft2 [57], Text2Video-Zero [46], ModelScope [47], replicating [3]’s implementation.
Performance is measured by overall classification accuracy and mean Average Precision (mAP). Table 2 summarizes the results from [3]. Baseline methods achieve moderate accuracies (45%–62%), with LNP [31] and Fusing [28] showing lower values. In contrast, our method obtains 97.06% average accuracy (Pika: 90.90%, VideoCraft2: 99.50%, Text2Video-Zero: 99.05%, ModelScope: 98.37%). Caveat: Comparing ReStraV to image-based detectors on a video task is not an even comparison (see next section for stronger baselines), yet it highlights the inadequacy of methods that rely solely on image-based features for AI-generated video detection neglecting temporal information.
@@
ReStraV (MLP) 90.90 99.50 99.05 98.37 97.06 99.12 98.76 98.93 98.44 98.81
7.2 ReStraV vs. video-based detectors.
@@
We compare ReStraV against the widely recognized VideoSwinTiny [40] (implementation from [17]) on the VidProM [3], with data setup following Section 7.1. Our evaluation considers three scenarios: Seen generators: Models are trained and tested on videos from a pool of the same four AI generators in Section 7.1, using a balanced set of 80,000 videos with a 50/50 train/test split. Unseen generators: Generalization is assessed by training models while excluding two specific AI generators (e.g., VC2 [59] and T2VZ [46]), which are then used for testing. Future generators: To simulate encountering a novel advanced model, ReStraV (trained on older generators) is tested on Sora [48]. Accuracy and mAP results are presented in Table 3.
To further test robustness, we evaluate ReStraV’s vs. nine SoTA video based detectors in Table 4. We consider two settings: Main (M) task, which is designed to test generalization across generators. Detectors are trained on videos from Pika [44], VideoCraft2 [57], Text2Video-Zero [46], ModelScope [47], and tested on MuseV [62], Stable Video Diffusion (SVD) [63], CogVideo [64], and Mora [65]. Plants (P) task, the most challenging subset from [17]. The challenge may arise from the complex and often stochastic nature (e.g., irregular leaf patterns, subtle wind movements), which can make generative artifacts less distinguishable from natural variations or harder for models to consistently detect (qualitative samples in Section 10.3). We use the same setting of task (M) but focusing on videos of plants in the test set. Baseline’s results from [17].
@@
(a) GenVidBench - Main (M) Task Acc. (%) Method MuseV SVD CogVideo Mora HD-VG Avg. [Nat.] TSM [34] 70.37 54.70 78.46 70.37 96.76 76.40 X3D [37] 92.39 37.27 65.72 49.60 97.51 77.09 MVIT V2 [38] 76.34 98.29 47.50 96.62 97.58 79.90 SlowFast [36] 12.25 12.68 38.34 45.93 93.63 41.66 I3D [35] 8.15 8.29 60.11 59.24 93.99 49.23 VideoSwin [39] 62.29 8.01 91.82 45.83 99.29 67.27 ReStraV (MLP) 93.52 94.01 93.52 92.97 91.07 93.01
(b) GenVidBench - Plants (P) Task Acc. (%) Method MuseV SVD CogVideo Mora HD-VG Avg. [Nat.] SlowFast [36] 81.63 29.80 75.31 19.31 73.03 55.30 I3D [35] 39.18 23.27 91.98 78.38 78.42 62.15 VideoSwin [39] 57.96 7.35 92.59 47.88 98.76 52.86 TPN [41] 43.67 20.00 85.80 86.87 94.61 64.24 UniFormer V2 [42] 13.88 7.76 41.98 95.75 97.93 64.76 TimeSformer [43] 77.96 29.80 96.30 93.44 87.14 75.09 ReStraV (MLP) 95.06 97.83 92.38 91.24 93.31 96.96
This robustness extends to the challenging Plants (P) task, where ReStraV obtains accuracies of $95.06\%$ (MuseV), $97.83\%$ (SVD), $92.38\%$ (CogVideo), $91.24\%$ (Mora), and $93.31\%$ (HD-VG/130M), leading to a $93.96\%$ average. The notable success on this task, particularly against generators like MuseV and SVD, may be attributed to ReStraV’s ability to capture specific curvature patterns inherent to “Plants” videos, which differ from more general generative artifacts. ReStraV remains highly effective and consistently competitive overall, as visualized by its position relative to the baseline spread (Fig. 12 in Section 10.4).
7.3 ReStraV vs. VLM detector on Physics-IQ dataset (matched real and generated videos).
@@
As a third and perhaps the most challenging test, we assess ReStraV on matched pairs of natural and generated videos from the Physics-IQ dataset. This dataset consists of real-world physical interactions and is special in the sense that it consists of both natural and AI-generated videos (198 per source) that are based on the very same starting frame(s): identical scenes, identical objects, identical lighting conditions as described in Section 5. We report [26]’s evaluation using a two-alternative forced-choice (2AFC) paradigm (a gold-standard psychophysical protocol). In each trial, a model sees a pair of videos: one real, one AI-generated.
The task of identifying the AI-generated video, is especially challenging due to the matched video nature. Motamed et al. [26] report results for a vision-language model (VLM), Gemini 1.5 Pro [61], on this dataset: Gemini identifies Runway and Pika videos with reasonable accuracy (74.8% and 80.5% respectively), but Sora videos prove challenging (55.6%, near 50% chance) due to their photorealism.
We evaluate ReStraV in the same setting, comparing against reported numbers [26]. We compute mean aggregated video curvature (Eq. 1) for each video and predict the one with higher mean curvature as “AI-generated.” No further classifier training or calibration is performed. Fig. 7 shows the results: ReStraV attains 97.5% for Pika [44], 94.9% for Runway [49], and 99.0% for Sora. This near-perfect performance across all three generators demonstrates that simple curvature statistics robustly discriminate real from generated videos without model fine-tuning.
{mdframed}
[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = NatBlue!20, roundcorner = 4pt, ] Takeaway 3: ReStraV demonstrates robust generalization across diverse generators and OOD scenarios, showing neural representation trajectories (distance $d$ and curvature $\theta$) as an effective paradigm for AI video detection.
@@ 8 Discussion
Our method is invariant to playing a video backwards. This is clearly unnatural, if things, e.g., fall up instead of down; though at the same time this also would not be an instance of an AI-generated video. The arrow of time [68, 69, 70] can be a strong signal, but our current metric is completely invariant to a reversal of time.
@@
Goodhart’s law states that “when a measure becomes a target, it ceases to be a good measure”. Likewise, in the context of fake video detection, it is conceivable that someone developing a video model could train it in a way that optimizes for deceiving detection measures. This concern generally applies to all public detection methods, including ours. As a possible mitigation strategy, it may be helpful to employ several detection methods in tandem, since it may be harder to game multiple metrics simultaneously without sacrificing video quality.
Furthermore, as video models become more and more capable of generating realistic, natural-looking videos, it is possible that future video models may not show the same statistical discrepancies between real and generated videos anymore, though this is hard to predict in advance.
@@
AI-generated video increasingly fuels fake news and disinformation [71, 72, 15]. ReStraV aims to positively impact this by enhancing content authentication. With AI-driven fraud like deepfake scams reportedly surging (e.g., a reported 2137% rise in financial sector attempts over three years [73]), efficient detection methods like ReStraV are important. Our approach helps combat these malicious uses, safeguarding information integrity and potentially offering AI safety insights by characterizing AI-generated temporal data.
However, deploying detection technologies like ReStraV faces an "arms race" with evolving generation methods (see Limitations; also [22]). Additionally, biases inherited from pre-trained encoders (e.g., DINOv2 [25]) may cause fairness issues across diverse content [74].
Mitigating these risks demands ongoing research, transparency about limitations, and using detectors primarily to aid human judgment. Key strategies include careful contextual deployment, rigorous bias auditing and debiasing efforts [74], promoting media literacy [15], and advancing complementary methods like robust content watermarking [20]. Moreover, to ensure that detection tools themselves do not infringe on user privacy when deployed at scale, they can be developed and trained using privacy-by-design principles such as Federated Learning [75, 76]. The importance of such measures is increasingly highlighted by policy trends, such as those related to content provenance and the EU AI Act [77], all vital for a trustworthy digital ecosystem.
@@ Acknowledgements
9 AI Video Generation
@@
10 Qualitative and Quantitative Analysis
10.1 Qualitative Examples of Perceptual Trajectories
To provide a more intuitive understanding of how curvature trajectories differ, Figure 8 presents several qualitative examples. Each plot shows the sequence of calculated curvature values ($\theta_{i}$) for a natural video (blue line) and a corresponding AI-generated video (red line) from the Sora [48], Pika [44], or Runway [49] models, after processing through the DINOv2 encoder as in Section 5.
@@ 10.1 Qualitative Examples of Perceptual Trajectories
10.2 ReStraV Classifiers Results Analysis
@@
Figure 9A presents the ROC curves for all classifiers detailed in Table 1. The curve for the MLP (ReStraV) is closest to the top-left corner and with the largest area, AUROC (98.63%). Figure 9B displays the normalized confusion matrices for the ReStraV classifiers from Section 6. The strong diagonal elements (e.g., correctly identifying natural videos as “Nat” and AI-generated as “GenAI”) and low off-diagonal values highlight the effectiveness. Specifically, correctly classifies a high percentage of both natural and AI-generated instances, aligning with the balanced accuracy and individual precision/recall/specificity metrics reported in Table 1.
Figure 10: Decision boundaries for Logistic Regression (LR) and Multi-Layer Perceptron (MLP) ResTraV’s classifiers. The plots illustrate how these models partition a 2D projection of the feature space (detailed in Section 6) to distinguish between natural (blue circles) and AI-generated (red triangles) videos. This comparison highlights the different decision boundaries learned by a linear (LR) and a non-linear (MLP) model when applied to the geometric trajectory features.
Fore demonstration purpose, we construct Voronoi tessellations to visualize the decision boundaries obtained from two different classification models: Logistic Regression (LR) and a Multi Layer Preceptor (MLP) classifier from Section 6 in Fig. 10. The visualization underscores the benefit of using a non-linear classifier for this task, given the nature of the feature space derived from ReStraV’s geometric analysis.
10.3 Qualitative Examples of “Plants” Task
Figure 11: Sample frames from the GenVidBench “Plants” task [17] for natural video (HD-VG/130M) and AI-generated plant video frames (MuseV [62], SVD [63], CogVideo [64], and Mora [65]).

### v2

@@ 7 Benchmark results
We evaluate ReStraV (with MLP from Section˜6) under four settings: A) Vs. SoTA image-based detectors on VidProM [3]; B) Vs. SoTA video detectors on VidProM [3] (evaluating “seen”, “unseen”, and “future” generator scenarios) and on the GenVidBench [17] dataset (in “cross-source”/“generator” (M) scenarios and its “Plants” hard task subclass (P)); C) Extreme generalization tests including one-to-many detection on the DeMamba [44] benchmark [44] and D) zero-shot evaluation on the Veo3 model [66]; E) Vs. a Vision-Language Model (VLM) performance (Gemini 1.5 Pro [67]) on the Physics‐IQ dataset [26] using matched real/generated video pairs. The best scores are bold, second best are underlined and the best method overall is highlighted in blue.
A) ReStraV vs. image-based detectors.
We evaluate our method ReStraV against eight SoTA image based detectors in Table˜2. ReStraV is trained with data and processing from Section˜6. We use a balanced test set (40,000 real videos; 10,000 AI-generated for each of four models: Pika [47], VideoCraft2 [68], Text2Video-Zero [49], ModelScope [50], replicating [3]’s implementation.
Performance is measured by overall classification accuracy and mean Average Precision (mAP). Table˜2 summarizes the results from [3]. Baseline methods achieve moderate accuracies (45%–62%), with LNP [31] and Fusing [28] showing lower values. In contrast, our method obtains 97.06% average accuracy (Pika: 90.90%, VideoCraft2: 99.50%, Text2Video-Zero: 99.05%, ModelScope: 98.37%). Caveat: Comparing ReStraV to image-based detectors on a video task is not an even comparison (see next section for stronger baselines), yet it highlights the inadequacy of methods that rely solely on image-based features for AI-generated video detection neglecting temporal information.
@@
ReStraV 90.90 99.50 99.05 98.37 97.06 99.12 98.76 98.93 98.44 98.81
B) ReStraV vs. video-based detectors.
We firstly compare ReStraV against the widely recognized VideoSwinTiny [40] (implementation from [17]) on the VidProM [3], with data setup following Section˜7. Our evaluation considers three scenarios: Seen generators: Models are trained and tested on videos from a pool of the same four AI generators in Section˜7, using a balanced set of 80,000 videos with a 50/50 train/test split. Unseen generators: Generalization is assessed by training models while excluding two specific AI generators (e.g., VC2 [63] and T2VZ [49]), which are then used for testing. Future generators: To simulate encountering a novel advanced model, ReStraV (trained on older generators) is tested on Sora [69]. Accuracy and mAP results in Table˜3.
@@
We futher evaluate ReStraV’s vs. nine SoTA video based detectors in Table˜4. We consider two settings: Main (M) task, which is designed to test generalization across generators. Detectors are trained on videos from Pika [47], VideoCraft2 [68], Text2Video-Zero [49], ModelScope [50], and tested on MuseV [70], Stable Video Diffusion (SVD) [71], CogVideo [72], and Mora [73]. Plants (P) task, the most challenging subset from [17]. The challenge may arise from the complex and often stochastic nature (e.g., irregular leaf patterns, subtle wind movements), which can make generative artifacts less distinguishable from natural variations or harder for models to consistently detect (qualitative samples in Section˜A.4). We use the same setting of task (M) but focusing on videos of plants in the test set. Baseline’s results from [17].
This robustness extends to the challenging Plants (P) task, where ReStraV obtains accuracies of $95.06\%$ (MuseV), $97.83\%$ (SVD), $92.38\%$ (CogVideo), $91.24\%$ (Mora), and $93.31\%$ (HD-VG/130M), leading to a $93.96\%$ average. This success may be attributed to ReStraV’s ability to capture specific curvature patterns inherent to “Plants” videos, which differ from more general artifacts. ReStraV remains highly effective as visualized by its position relative to the baseline spread (Fig.˜13 in Section˜A.5).
@@
(a) GenVidBench - Main (M) Task Acc. (%)
Method MuseV SVD CogVideo Mora HD-VG Avg.
[Nat.]
TSM [34] 70.37 54.70 78.46 70.37 96.76 76.40
X3D [37] 92.39 37.27 65.72 49.60 97.51 77.09
MVIT V2 [38] 76.34 98.29 47.50 96.62 97.58 79.90
SlowFast [36] 12.25 12.68 38.34 45.93 93.63 41.66
I3D [35] 8.15 8.29 60.11 59.24 93.99 49.23
VideoSwin [39] 62.29 8.01 91.82 45.83 99.29 67.27
ReStraV 93.52 94.01 93.52 92.97 91.07 93.01
(b) GenVidBench - Plants (P) Task Acc. (%)
Method MuseV SVD CogVideo Mora HD-VG Avg.
[Nat.]
SlowFast [36] 81.63 29.80 75.31 19.31 73.03 55.30
I3D [35] 39.18 23.27 91.98 78.38 78.42 62.15
VideoSwin [39] 57.96 7.35 92.59 47.88 98.76 52.86
TPN [41] 43.67 20.00 85.80 86.87 94.61 64.24
UniFormer V2 [42] 13.88 7.76 41.98 95.75 97.93 64.76
TimeSformer [43] 77.96 29.80 96.30 93.44 87.14 75.09
ReStraV 95.06 97.83 92.38 91.24 93.31 96.96
C) One-to-many generalization Test.
We reproduced the one-to-many task from DeMamba [44], training on a single generator and testing on multiple unseen ones (Sora [69], MorphStudio [74], Gen2 [75], HotShot [76], Lavie [5], Show [77], MoonValley [78], Crafter [68], ModelScope [50] and WildScrape[44]). Table 5 shows average results across three training conditions. Notably, ReStraV achieves competitive or superior scores in most scenarios against specialized video detectors (TALL [40], NPR [79], STIL [80], and DeMamba [44]) using only the extracted trajectories.
Table 5: One-to-many: training on one generator, testing on unseen generators (Sora, MorphStudio, Gen2, HotShot, Lavie, Show-1, MoonValley, Crafter, ModelScope, WildScrape). Avarage results from DeMamba benchmark [44]. $\uparrow$ is better.
Train: Pika Train: SEINE Train: OpenSora
Method R F1 AP R F1 AP R F1 AP
NPR 0.514 0.531 0.650 0.462 0.539 0.611 0.593 0.523 0.576
STIL 0.738 0.517 0.630 0.724 0.506 0.608 0.434 0.489 0.526
TALL 0.714 0.557 0.623 0.657 0.609 0.681 0.492 0.532 0.571
DeMamba 0.757 0.726 0.817 0.810 0.787 0.894 0.738 0.671 0.738
ReStraV 0.735 0.827 0.797 0.820 0.898 0.854 0.771 0.797 0.717
D) Zero-shot Generalization Test.
We tested zero-shot generalization on Google’s Veo3 [66], a state-of-the-art model acclaimed for its ability to generate videos with plausible physical interations and consistent object interactions (qualitative frame samples in Section˜A.6). Using only the MLP trained in Section 6 (without any Veo3 videos in training), we tested on 200 Veo3 versus 200 natural video pairs and achieved 83.2% accuracy, 85.1% F1, and 86.9% AUROC. This shows that the curvature-based detection signal can generalizes to future generators not represented in the training distribution, supporting our hypothesis that current generative models fundamentally struggle to replicate the temporal smoothness characteristic of natural world dynamics in learned representation spaces.
E) ReStraV vs. VLM detector on Physics-IQ dataset (matched real and generated videos).
@@
As a third and perhaps the most challenging test, we assess ReStraV on matched pairs of natural and generated videos from the Physics-IQ dataset. This dataset consists of real-world physical interactions and is special in the sense that it consists of both natural and AI-generated videos (198 per source) that are based on the very same starting frame(s): identical scenes, identical objects, identical lighting conditions as described in Section˜5. We report [26]’s evaluation using a two-alternative forced-choice (2AFC) paradigm (a gold-standard psychophysical protocol). In each trial, a model sees a pair of videos: one real, one AI-generated.
The task of identifying the AI-generated video is especially challenging due to the matched video nature. Motamed et al. [26] report results for a VLM, Gemini 1.5 Pro [67]. Gemini identifies Runway and Pika videos with reasonable accuracy (74.8% and 80.5% respectively), but Sora videos prove challenging (55.6%, near 50% chance) due to their photorealism.
We evaluate ReStraV in the same setting, comparing against reported numbers [26]. We compute mean aggregated video curvature (Eq.˜1) for each video and predict the one with higher mean curvature as “AI-generated.” No further classifier training or calibration is performed. Fig.˜7 shows the results: ReStraV attains 97.5% for Pika [47], 94.9% for Runway [81], and 99.0% for Sora. This near-perfect performance across all three generators demonstrates that simple curvature statistics robustly discriminate real from generated videos without model fine-tuning.
Takeaway 3: ReStraV demonstrates robust generalization across diverse generators and OOD scenarios, showing neural representation trajectories (distance $d$ and curvature $\theta$) as an effective paradigm for AI video detection.
@@ 8 Discussion
Our method is invariant to playing a video backwards. This is clearly unnatural, if things, e.g., fall up instead of down; though at the same time this also would not be an instance of an AI-generated video. The arrow of time [84, 85, 86] can be a strong signal, but our metric is invariant to a reversal of time.
@@
Goodhart’s law states that “when a measure becomes a target, it ceases to be a good measure”. Likewise, in the context of fake video detection, it is conceivable that someone developing a video model could train it in a way that optimizes for deceiving detection measures. This concern generally applies to all public detection methods, including ours. As a possible mitigation strategy, it may be helpful to employ several detection methods in tandem, since it may be harder to game multiple metrics simultaneously without sacrificing video quality. Furthermore, as video models become more and more capable of generating realistic, natural-looking videos, it is possible that future video models may not show the same statistical discrepancies between real and generated videos anymore, though this is hard to predict in advance.
@@
AI-generated video increasingly fuels fake news and disinformation [87, 88, 15]. ReStraV aims to positively impact this by enhancing content authentication. With AI-driven fraud like deepfake scams reportedly surging (e.g., a reported 2137% rise in financial sector attempts over three years [89]), efficient detection methods like ReStraV are becoming fundamental.
However, deploying detection technologies like ReStraV faces an “arms race” with evolving generation methods (see Limitations; also [22]). Additionally, biases inherited from pre-trained encoders (e.g., DINOv2 [25]) may cause fairness issues across diverse content [90]. Mitigating these risks demands ongoing research, transparency about limitations, and using detectors primarily to aid human judgment. Key strategies include careful contextual deployment, rigorous bias auditing and debiasing efforts [90], promoting media literacy [15], and advancing complementary methods like robust content watermarking [20].
The importance of AI-safety measures [91] is increasingly reflected in policy initiatives like the Coalition for Content Provenance and Authenticity [92] and the EU AI Act [93], both vital for a trustworthy digital ecosystem. However, deploying detection tools at scale presents its own challenges, especially concerning user privacy. To address this, detectors can be distributed and trained using privacy-by-design principles [94, 95, 96]. Within this framework, tools like ReStraV are crucial for ensuring the digital ecosystem remains grounded in reality, providing a critical defense against the long-term risk of epistemic decay in world models [97, 98].
@@ Acknowledgements
B.1 Impact of Video Length and Sampling Density ........................................................................................................................................................................ B.1
B.2 Robustness to Temporal Window Position ........................................................................................................................................................................ B.2
Appendix A AI Video Generation
@@
A.1 Qualitative Examples of Perceptual Trajectories
To provide a more intuitive understanding of how curvature trajectories differ, Figure 8 presents several qualitative examples. Each plot shows the sequence of calculated curvature values ($\theta_{i}$) for a natural video (blue line) and a corresponding AI-generated video (red line) from the Sora [69], Pika [47], or Runway [81] models, after processing through the DINOv2 encoder as in Section˜5.
@@ A.1 Qualitative Examples of Perceptual Trajectories
A.2 ReStraV Classifiers Results Analysis
@@
Figure 9A presents the ROC curves for all classifiers detailed in Table 6. The curve for the MLP (ReStraV) is closest to the top-left corner and with the largest area, AUROC (98.63%). Figure 9B displays the normalized confusion matrices for the ReStraV classifiers from Section˜6. The strong diagonal elements (e.g., correctly identifying natural videos as “Nat” and AI-generated as “GenAI”) and low off-diagonal values highlight the effectiveness. Specifically, correctly classifies a high percentage of both natural and AI-generated instances, aligning with the balanced accuracy and individual precision/recall/specificity metrics reported in Table 6.
Figure 10: Decision boundaries for Logistic Regression (LR) and Multi-Layer Perceptron (MLP) ResTraV’s classifiers. The plots illustrate how these models partition a 2D projection of the feature space (detailed in Section˜6) to distinguish between natural (blue circles) and AI-generated (red triangles) videos. This comparison highlights the different decision boundaries learned by a linear (LR) and a non-linear (MLP) model when applied to the geometric trajectory features.
For demonstration purpose, we construct Voronoi tessellations to visualize the decision boundaries obtained from two different classification models: Logistic Regression (LR) and a Multi Layer Preceptor (MLP) classifier from Section˜6 in Fig.˜10. The visualization underscores the benefit of using a non-linear classifier for this task, given the nature of the feature space derived from ReStraV’s geometric analysis.
A.3 Feature Importance Analysis
Figure 11: Permutation Feature Importance.The plot shows the mean drop in F1-score (%) when each feature is permuted, with error bars indicating standard deviation. Features are grouped into “Aggregated Statistics” (e.g., mean, variance) and “Individual Time-Step Signals” (e.g., $d_{0}$, $\theta_{1}$). The results highlight the overwhelming importance of the mean curvature ($\mu_{\theta}$) and other aggregated statistics in distinguishing natural from AI-generated videos.
To identify which of the 21 features contributed most to the performance of our best classifier (the MLP), we conducted a permutation feature importance analysis. We trained the MLP on the 50,000 AI-generated and 50,000 natural videos described in Section 5 and then measured the drop in F1-score when each feature was individually shuffled. A larger drop indicates a more important feature. The results are visualized in Fig.˜11.
The mean curvature ($\mu_{\theta}$) is unequivocally the most critical feature. Its importance is more than double that of the next most influential feature, the curvature variance ($\sigma^{2}_{\theta}$). This provides strong quantitative evidence that the overall “straightness” of a trajectory is the primary signal our method leverages. The eight aggregated statistical features occupy the top tiers of importance. In contrast, the features representing individual, time-specific distance and curvature values ($d_{i},\theta_{i}$) have a much smaller impact, suggesting they provide complementary but less critical information. These findings validate the distributions shown in Figure˜5, where the aggregated statistics for curvature and distance showed the clearest separation between natural and AI-generated videos. The classifier effectively learns to exploit these high-level geometric properties.
A.4 Qualitative Examples of “Plants” Task
Figure 12: Sample from the GenVidBench “Plants” task [17] for natural video (HD-VG/130M) and AI-generated plant video frames (MuseV [70], SVD [71], CogVideo [72], and Mora [73]).